Subject: 3.5-m Users Committee minutes 7/17/06
From: Bruce Gillespie
Submitted: Thu, 20 Jul 2006 16:08:52 -0600
Message number: 979
(previous: 978,
next: 980
up: Index)
APO 3.5-m Users Committee Phonecon, 7/17/06
Attending: Suzanne Hawley, Bruce Gillespie, Michael Strauss, Rene
Walterbos, John Bally, Remy Indebetouw, Al Harper, Scott Anderson,
Jon Fulbright, John Wilson, Russet McMillan, Anjum Mukadam, Russell Owen
Absent: Jon Holtzman, Bruce Balick
Minutes taken by Bruce Gillespie
**********************************
User feedback, comments from institutional representatives:
Princeton (Michael Strauss): Nothing much to report, everything is
fine.
Colorado (John Bally): No issues.
NMSU (Walterbos): Nothing to report.
Johns Hopkins (Jon Fulbright): Nothing to report.
Chicago (Al Harper): Nothing new.
UWashington (Scott Anderson & Suzanne Hawley): No issues or problems.
Some UW staff are working on Agile (see below). Also, grad students
are learning to use NIC-FPS.
Univ. of Virginia (Remy Indebetouw): Nothing to report.
**********************************
Discussion of telescope/instruments report:
There was discussion about the new DIS medium gratings (see message
978 in the apo35-general archive from Stauss/McMillan). The gratings
appear to be very good and will hereafter be the default medium-
dispersion gratings. Also, the newly installed altitude and azimuth
drive boxes seem to have made an improvement in the telescope
tracking performance. Otherwise, Bruce gave a brief overview of the
report, given below. Russet added that Gabrelle Saurage, our new
Observing Specialist, has completed training and has soloed on the
telescope.
*************************
3.5-m Telescope, Instruments, and CIF Projects Highlights,
6/09/06 through 7/12/06
Bruce Gillespie
0) Overview
We were happy to greet the early onset of the monsoonal rains, which
dumped a few inches at the site, and to our great relief has greatly
reduced the forest fire risk. There were periods of decent observing
conditions, two new gratings were installed in DIS, the new altitude
drive box was successfully installed, and an unexpected problem with
the tertiary mirror mount occurred and was repaired with little
effective lost telescope time.
1) Telescope
The new version of TUI saw extensive use, seems to work well, and
minor bugs were found and quickly fixed. During a short planned
engineering downtime, the new altitude drive box was installed and
tested. All three axis drive motors (one altitude, two azimuth) are
now coupled to the newly fabricated boxes, which seem to have
improved telescope tracking and pointing performance, as well as
having given us a better/faster maintenance path if there are drive
box problems in the future.
The biggest problem during the period was the delaminating of the
metal center support post from the back of the tertiary mirror. This
was an unexpected reprise of a similar problem we had last fall, and
after careful engineering analysis and discussions with a materials
expert at the Steward Mirror Lab, it was determined that the metal
puck that is attached to the glass was probably too small in contact
area, and also that the RTV adhesive has marginal performance in this
particular application. A larger puck was machined, and a fresh
supply of adhesive was used to reattach the support post. The
telescope was returned to on-sky checkout and science five days after
the problem was diagnosed; nearly all of the lost observing time was
unusable for science anyway due to clouds and rain.
2) Instruments
CorMASS was used with good results for the first time in a while.
NIC-FPS focus scripts were tested with mixed results; more tests are
planned. There are unconfirmed reports that the "haze" in DIS blue
camera has returned, which will be addressed shortly. SPIcam was
given its annual vacuum pumping. The new blue and red medium
resolution DIS gratings were received, installed, aligned,
characterized, and accepted. The NIC-FPS instrument team attempted
to commission the Fabry-Perot etalon, and had difficulties (which
John Bally will report on).
3) Engineering and CIF projects
Several of the engineering items are reported above and not repeated
here. In addition to the two new DIS gratings, the new red
replacement CCD has been accepted and is being shipped to APO. Its
installation, plus that of the new red field flattener and new blue
prism are now being planned for September. The top-end project is
going well and is scheduled for installation on the telescope
beginning 28 August.
4) Miscellaneous
Fire risk in the Lincoln Nat'l forest is now low and the forest has
been reopened to the public.
**********************************
DIS upgrade status - Hawley:
Suzanne gave a summary report on the DIS upgrade items. The two
medium-dispersion gratings have been installed and accepted, and are
now the default. She would appreciate hearing feedback from DIS
users on how well they are working. The new red CCD has been
accepted and we expect delivery at APO shortly. The blue prism is
expected by the end of July. Due to manpower constraints, we are now
planning to install the new CCD, red field flattener, and blue prism
in early September. A new high-dispersion grating is also expected
to be delivered and installed in the next month or so.
Michael Strauss noted that we have a long-standing policy that all
3.5-m users be signed up on the apo35general mailer, and he has
reasons to think that many new users are not getting our
announcements (e.g., about the DIS new gratings).
ACTION: Users Committee members should send Michael
(strauss@astro.princeton.edu) the names and e-mail addresses of 3.5-m
users who are not yet signed up for the apo35general mailer.
**********************************
NIC-FPS Fabry-Perot status - Bally:
John reported that he, Stephane Beland, and Nathaniel Cunningham
recently were at APO to work on getting the NIC-FPS etalon
commissioned. They were unable to get the etalon controller to
behave in a stable fashion with the etalon at cryo operating
temperatures, and one of the actuator stacks appears to have become
broken during the tests, rendering the etalon out of order.
Conversations with the vendor's engineers did not lead to a
definitive conclusion as to what the problem was. Subsequent
discussion at CU led to several options to consider: 1) fix the
etalon again (it was also broken last year), 2) add an optical
monitoring port to the instrument to sense etalon stability, 3)
replace the etalon with a room-temperature Fabry-Perot that would be
mounted in front of NIC-FPS, and 4) use a tiltable narrow-band filter
instead of a Fabry-Perot. Further talks are planned to determine
which path(s) to take. John wrote a trip report detailing the
situation; copies can be e-mailed to interested persons by contacting
John (bally@casa.colorado.edu) or Bruce (gillespi@apo.nmsu.edu).
**********************************
"Agile" high-speed photometer - Hawley, Anjum Mukadum:
Anjum Mukadum and Paula Szkody have been awarded a grant to build a
high-speed CCD photometer to use on the 3.5-m telescope. It would be
used for variable white dwarf stars, CVs, flare stars, etc. The
instrument, called Agile, was brought to the attention of the Users
Committee last year, and is now a funded project. Anjum wrote a four-
page white paper describing the science goals and technical
specification of Agile, which was given to the Users Committee prior
to this meeting. Copies of this paper can be furnished to interested
persons by contacting Anjum (anjum@astro.washington.edu) or Bruce
(gillespi@apo.nmsu.edu). We are considering ways of implementing
Agile on the telescope that would enable it to be a fully integrated
facility instrument, available to all 3.5-m users through the TUI
interface. Michael asked why Agile needed so large a field of view,
and Anjum replied that the 2.6 x 2.6 arcmin field is necessary in
order to have comparison stars. She expects to have the instrument
ready for the telescope in six to eight months, and she invites
questions from potential users about its features.
**********************************
ACCORD System Workshop - Hawley:
Suzanne announced that AURA is planning to hold a 3rd "System
Workshop" and that ACCORD, which is comprised of the directors of US
observatories with telescopes 3m and larger, has been asked to take
an active role in the planning for the workshop which is planned for
mid November. Suzanne, Don York and others from APO participated in
the 2nd system workshop which was held in 2004. To promote active
participation from all levels of the US community, the ACCORD
directors were asked to furnish the names of science working group or
users group members to the organizing committee. Thus, Suzanne sent
the names of the users committee to the organizers, and they can
expect to perhaps see some email traffic about this meeting soon.
Other interested ARC members should feel free to contact Suzanne for
more information or to be added to the list furnished to the
organizing committee.
**********************************
TUI upgrade policy - Russell Owen:
Russell Owen presented the proposed TUI upgrade policy, which has
been under discussion for a few months. The policy was endorsed by
the committee members without modification, and is given for
reference below. Russell will consider embedding the policy inside
TUI itself, so that it is readily available to the end users.
**********************************
Comings and goings of committee members - Hawley:
Suzanne announced that Bruce Balick and Karl Glazebrook are retiring
from the Users Committee, and thanked them for many years of
service. Scott Anderson and Jon Fulbright are the new committee reps
from UW and JHU, respectively---welcome aboard, Scott and Jon.
Suzanne also mentioned that Bruce Balick has stepped down as
astronomy department chair at UW, and that she has assumed that role
as of this month.
**********************************
Previous Meetings Action Items report:
[open from previous months]:
ACTION: BruceG will ask Russell what the pros and cons are of
keeping TUI compatible with older versions of operating systems and
libraries. STATUS: closed (see policy below).
ACTION: BruceG to discuss issues of computer and network security
with Fritz Stauffer. STATUS: Working group (site staff and users)
formed to frame requirements.
ACTION: BruceG will ensure that the orientation checklist is up to
date and available on-line, so that prospective new users can review
it prior to their orientation site visit. STATUS: Closed, see
http://www.apo.nmsu.edu/35m_operations/35m_manual/General_Information/
Training.html.
ACTION: Jon Holtzman to talk to Fritz and BruceG about installing a
RAID backup at NMSU. STATUS: Closed, system up and running at NMSU.
[new actions from last month]:
ACTION: Suzanne asked each Users Committee member to send her a
paragraph or two in the next two days with an expression of interest
and/or questions regarding this proposed NIR instrument.
STATUS: Closed.
ACTION: BruceG to forward Futures documents to Users Committee
members, who then help the institutional Futures Committee member
share them with their faculties and gather feedback.
STATUS: Closed.
**********************************
Next meeting:
There will be no meeting next month. The next Users Committee
phonecon will be on Monday, 18 September, at 8:30 AM Pacific Time.
Agenda and other materials will be sent to the committee members
during the preceding week.
**********************************
Proposed TUI Upgrade Policy
Russell Owen
06/19/06
This proposal describes what APO expects from users as far as keeping
TUI current and the resources needed to run TUI. The intent is to
balance the needs of the community against the limited resources of APO.
** Which TUI Versions You May Use **
Obviously we can't insist that users upgrade TUI every few months.
It's too much work for the institutions. On the other hand, older
versions of TUI can be a problem for APO because the observing
specialists may find it more difficult to help the users and because
of the presence of known bugs.
Thus use of older versions of TUI will be permitted where possible,
but discouraged.
APO may occasionally require an update for various reasons, including:
- Incompatible changes to the "hub" software that are deemed
sufficiently important to justify loss of backwards compatibility.
- Known serious problems with a version of TUI.
- Protecting the telescope or instruments.
** The Resources You Need To Run TUI **
The details depend on whether TUI is a binary distribution (Mac or
Windows) or a source distribution (unix).
For binary distributions (Mac and Windows), the main concern is that
TUI may not be compatible with every version of the operating system.
I will try to support the two most recent major releases of the
operating system.
Source distributions (unix) are usually independent the operating
system, but require python, tcl/tk and some support libraries. I will
support the two most recent major versions of python. I will try to
minimize the number of support libraries and avoid changes to those
support libraries.
In addition, TUI uses a few optional standalone external packages
(xpa and ds9, plus possibly xephem in the future). TUI will support
the two most recent major releases of these packages and/or versions
that are approximately a year old.
Other comments:
- Unix: it would be handy to have a binary distribution (or at least
a more self-contained source distribution). Is there somebody who
could work on this? I haven't found the time to learn how to do it.
- Intel Macs: at present TUI runs in emulation mode. That is not
likely to change until it becomes a problem or making universal
binaries becomes easier.
- Windows: support is strictly on a "best effort" basis. I don't own
a Windows machine and have little knowledge of the operating system
and limited ability to test releases.
APO APO APO APO APO Apache Point Observatory 3.5m APO APO APO
APO
APO This is message 979 in the apo35-general archive. You can find
APO the archive on http://www.astro.princeton.edu/APO/apo35-general/INDEX.html
APO To join/leave the list, send mail to apo35-request@astro.princeton.edu
APO To post a message, mail it to apo35-general@astro.princeton.edu
APO
APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO APO